Check whether or not the sequence (n- 1/n) is convergent

1
Expert's answer
2022-06-06T16:43:26-0400
"\lim\limits_{n\to \infin}a_n=\lim\limits_{n\to \infin}(n-\dfrac{1}{n})"

"=\lim\limits_{n\to \infin}n-\lim\limits_{n\to \infin}\dfrac{1}{n}=\lim\limits_{n\to \infin}n-0=\infin"

The sequence "(n-\dfrac{1}{n})" is not convergent.